When solutions of silver nitrate and magnesium chloride are mixed, silver chloride precipitates out of solution according to the equation:

2AgNO3(aq) + MgCl2(aq) ---> 2AgCl(s) + Mg(NO3)2(aq)

What mass of silver chloride can be produced from 1.04L of a 0.195M solution of silver nitrate?
Answer #1

0.195 M AgNO3 means-

1 L solution contains 0.195 mole AgNO3.

Hence, 1.04 L solution contains (0.195/1)×1.04 mole = 0.203 mole AgNO3.

From the reaction stoichiometry we can say that,

0.203 mole AgNO3 will give 0.203 mole AgCl.

Molar mass of AgCl = 143.32 g/mole.

Mass in g = no. of moles×molar mass.

Hence, mass of AgCl produced = 0.203×143.32 g = 33 g.